OpenStudy (anonymous):

11. x (years since 1991) 2 5 7 8 y (book sales billions $) 18 21 23 24 Using the equation for this table and linear extrapolation, predict the book sales, in billions of dollars, in 2010, using data from the previous question. (Note that x is the number of years after 1991.) (Points : 1) 35 billion 48 billion 72 billion 12. Write an equation in slope-intercept form of the line with an x-intercept of 6 and a y-intercept of 5.(Points : 1) y = 6x + 5 y = -5/6 x + 5 y = -6x + 5 y = 5/6 x + 5

OpenStudy (blurbendy):

y = x + 16 2010 is 19 years after 1991 y = 19 + 16 y = 35 billion

OpenStudy (blurbendy):

(6, 0) (0,5) (5 - 0) = 5 (0 -6) = -6 y = -5/6x + 5